OpenStudy (anonymous):

Evaluate the function f(x) = x² + 1 for f(2). a.1 b.5 c.7

OpenStudy (anonymous):

b

OpenStudy (callisto):

f(2) = 2² + 1 = 4+1 = 5=b

OpenStudy (anonymous):

B. F(2) = 2^2 + 1 = 4+1 = 5

OpenStudy (anonymous):

5

OpenStudy (anonymous):

5